Solve the system
4y=12
X+4y+z=15
y-3z=0

Answer by Alan3354(69443)

Solve the system
4y=12 --> y = 3
X+4y+z=15
y-3z=0 y = 3, find z
------------
X+4y+z=15 y & z are known, find x